Circular No. 7410

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000cb IN IC 1158 M. Papenkova and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 7406), report the discovery, on an unfiltered image taken on Apr. 27.4 UT with the 0.8m Katzman Automatic Imaging Telescope (KAIT), of an apparent supernova (mag ∼ 18.0) located at α = 16^{h}01^{m}32^{s}.15, δ = +1^{o}42′23′′.0 (equinox 2000.0), which is 29′′.5 west and 4′′.2 south of the nucleus of IC 1158. A KAIT image taken on Apr. 24.4 also shows SN 2000cb at mag ∼ 18.5, but an Apr. 9.4 image shows nothing at this position (limiting mag ∼ 19.0).

S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2000cb, taken by P. Berlind on Apr. 28.4 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova at a very early epoch. The spectrum exhibits a slightly blue continuum with a prominent P-Cyg Hα profile, as well as broad absorption features of Hβ and Hγ. Adopting the NASA/IPAC Extragalactic Database recession velocity of 1927 km/s for the host galaxy yields high expansion velocities for the supernova: 18 300 (from the Hα profile), 15 900 (Hβ), and 14 400 (Hγ) km/s.

G. Aldering and A. Conley, Lawrence Berkeley National Laboratory, report that an optical spectrum of SN 2000cb, taken with the Cerro Tololo 4-m telescope (+ RC spectrograph) on Apr. 29.3 UT, reveals that the object to be a type-II supernova near maximum. The spectrum shows a well developed Hα P-Cyg profile, Hβ absorption, and weak He I at 587.5 nm. The location of the Hα absorption minimum, along with the host recessional velocity of 1927 km/s (given by NED), yields a photospheric expansion velocity of 10 000 km/s. The lack of an apparent narrow Na D feature suggests little extinction by the host galaxy.

COMET C/2000 H1 (LINEAR) An object reported by LINEAR (discovery observation below) and

posted on the NEO Confirmation Page has been noted to be diffuse at Kleť (Apr. 27.0 UT), Ondřejov (Apr. 27.1), and Modra (Apr. 29.0). M. Hicks (Table Mountain Observatory) reported a 7′′ central condensation and 20′′ tail in p.a. ∼ 120^{o} on Apr. 29.4. Full astrometry and preliminary parabolic elements appear on MPEC 2000-H40.

2000 UT  α_{2000}  δ_{2000}  m_{2}
Apr. 26.38774  18^{h}28^{m}39.^{s}85  +47^{o}18′35′′.9  18.7
2000 April 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green